### v1

@@ 2.2.2 BRaTS dataset
In our second stage, we perform self-supervised pretraining on MRI images from the training set of BraTS 2021 Task 1 (Tumor Segmentation) challenge. This dataset consists of 1,251 subjects each with T1, T1-contrast enhanced (T1-ce), T2, and T2-FLAIR images. We obtained all publicly available images as part of the challenge. The BraTS challenge utilizes a similar standard preprocessing pipeline as the UKB dataset. First, images are converted from DICOM images to NIfTI using dcm2niix ([li_first_2016]) and tools available from the Cancer Imaging Phenomics Toolkit (CaPTk) ([davatzikos_cancer_2018]). Images are then co-registered to the SRI24 template and resampled to a uniform resolution of 1 mm3. Finally, each modality is skull stripped; through skull stripping and conversion to NIfTI, images are both defaced, and all identifiable information is stripped from the file metadata ([baid_rsna-asnr-miccai_2021]). Imaging volumes were then segmented into three tumor classes using the STAPLE algorithm ([warfield_simultaneous_2004]) across previous BraTS winners and refined manually. These manual annotations were further verified by multiple board-certified neuro-radiologists, resulting in quality-controlled tumor segmentation labels across all 4 modalities in 3 classes: Gd-enhancing tumor (referred to as the whole tumor: WT), edematous tissue (ED), and necrotic tumor core (TC).
@@
Additionally, we perform self-supervised Stage 2 pretraining and fine-tuning on MRI images from the training set of the Anatomical Tracings of Lesions After Stroke (ATLAS) v2.0 Dataset [liew_large_2022]. This dataset consists of 655 T1-ce MRIs aggregated from 44 research cohorts. Each MRI is from one subject, and time-points range from <24 hours to >180 days after stroke onset. The standard labeling pipeline for the ATLAS dataset consists of (1) manual quality control to exclude large motion artifacts, (2) manual lesion tracing in ITK-SNAP [yushkevich_user-guided_2006] [yushkevich_itk-snap_2017], and (3) mask review by two independent raters. The data then went through a similar prepossessing pipeline to BraTS; MR images were first intensity-normalized with the MINC toolkit (https://github.com/BIC-MNI/minc-toolkit). The MINC toolkit was also used to then register the images to the MNI-152 template. Finally, FreeSurfer’s MRI deface functionality was used to deface the scans. Images were reviewed again in a final quality check at the end of the pipeline before being included in the dataset. Segmentations are evaluated on four metrics - Dice coefficient for the final segmentation (Dice), the difference between true total lesion volume and predicted total lesion volume (Volume Difference), the difference in the number of lesions between ground truth and prediction (Lesion Count), and Lesion-wise F1 Score. The Lesion-wise F1 Score is calculated by performing 3D connected-component analysis to determine true positives, false positives, and false negatives. A true positive is defined as any 3D connected-component in the ground-truth image that overlaps with at least one voxel in the prediction image. Conversely, a false positive is identified as any 3D connected-component in the prediction image that does not overlap with the ground-truth image, while a false negative is a connected-component in the ground-truth that lacks any overlapping voxels in the prediction image [liew_large_2022].
@@ 2.3 Stage 1: Pretraining on the UKB
The initial stage of pretraining involves the self-supervised learning of a trans-former-based neural network model using a substantial unlabeled image dataset. For this purpose, the UKB dataset ([littlejohns_uk_2020]) is utilized. From our 82,800 3-dimen-sional volumetric images used for pretraining, the input MRI modalities are randomly cropped into 96x96x96 sub-volumes and augmented with random inner cutout and rotation. These augmented images are then fed into the SwinUNETR encoder for processing.
Adopting the methodology from [tang_self-supervised_2022], the SwinUNETR encoder is pretrained through three distinct proxy tasks that serve as self-supervised fine-tuning mechanisms. The SwinUNETR architecture incorporates a Swin Transformer encoder, which is specifically designed to handle 3D input patches. This encoder operates with a patch size of 2x2x2, a feature dimension of 8, and an embedding space of 48 dimensions. It consists of four stages, with a patch merging layer introduced between stages to reduce the feature size by half.
The primary objective of pretraining is to minimize the total loss function. This is achieved by training the SwinUNETR encoder using three proxy tasks: masked volume inpainting, 3D image rotation, and contrastive coding. To cater to varying complexities, three models have been developed: the foundational BrainFounder-Tiny with 62 million parameters, the intermediate BrainFounder-Small with 64 million parameters, and the advanced BrainFounder-Big, which boasts 69 million parameters. The key differentiation among these models is the variation in the number of sliding window blocks within their third stage. BrainFounder’s sliding-window encoder backbone’s parameters, number of SSL heads, and number of sliding window blocks are shown in Table 2.
For the pretraining process, a total of 64 NVIDIA DGX A100 GPUs, distributed across 8 DGX-2 nodes, are deployed at the University of Florida’s HiPerGator-AI supercomputer. Data parallelism is implemented to optimize the efficiency of model training. Both training and validation losses are monitored to track progress. The AdamW optimizer is employed with a warm-up cosine scheduler set for 500 iterations. The training employs a batch size of 2 per GPU, using 96x96x96 patches. The initial learning rate is established at 6E-6, coupled with a momentum of 0.9 and a decay of 0.1 over 15,000 iterations. These parameters are summarized in Table 3.
BrainFounder-Tiny (62M) BrainFounder-Small (64M) BrainFounder-Big (69M) # of encoder parameters 19,097,191 20,982,103 26,636,839 Encoder layer level Output size # of SSL Heads # Swin Blocks # of SSL Heads # Swin Blocks # of SSL Heads # Swin Blocks Level 1 48x(48x48x48) 3 2 3 2 3 2 Level 2 96x(24x24x24) 6 2 6 2 6 2 Level 3 192x(12x12x12) 12 2 12 6 12 18 Level 4 384x(6x6x6) 24 2 24 2 24 2 Feature size 48 48 48 Bottleneck dimension 768 768 768
@@
Fig. 4: Training (left) and validation (right) loss of Stage 1-pretraining three different scale of BrainFounder models on UKB.
@@ 2.4.1 Stage 2: Pretraining on BraTS
Given that the BraTS dataset comprises four modalities, but only two (specifically, T1 and T2) were selected for pretraining in the initial stage, the first layer of the pretrained network on UKB was modified. This modification involved expanding the number of input channels by adding two new channels, whose weights were randomly initialized using the Kaiming initialization method ([he_delving_2015]).
Hyperparameter settings for Stage-2 Pretraining can be found in Table 3. For pretraining on BraTS, two NVIDIA A100 GPUs, each with 32 GB of memory, were utilized. Depending on the model size, the BrainFounder models require between 48 to 72 hours for training. The batch size and learning rate were uniformly set at 2 and $1\times 10^{-4}$ for all models during this pretraining phase.
2.4.2 Fine-tuning on BraTS
@@ 2.4.3 Few-shot Learning on BraTS
To investigate our model’s performance using limited training data, we conducted a systematic comparison between BrainFounder and the baseline model, SwinUNETR, utilizing a descending percentage training approach in the context of the BraTS challenge. Using both our BrainFounder pretrained model and SwinUNETR, we finetuned on 100% of the BraTS training dataset, with subsequent incremental reductions in data availability, decreasing to a final 5% of the original dataset. This method aimed to explore the impact of training data scarcity on model performance and adaptability. Performance evaluations were carried out on the BraTS test set after each training step and evaluated with the Dice coefficient to assess segmentation accuracy.
@@
Stage Data No. Subjects GPU Batch size Learning rate No. steps No. input channel Encoder only Pretraining UKB 43369 64 x A100 128 6E-06 2E+05 2 Encoder + Decoder Pretraining & Fine-tuning BRaTS 1251 2 x A100 2 1E-04 50000 4 Encoder + Decoder Pretraining & Fine-tuning ATLAS v2.0 655 4 x A100 4 5E-03 600 1
2.4.4 Modality Restriction and Flexibility of Modalities in Training and Inference
The proposed method is designed to be adaptable across various data modalities in downstream tasks. Specifically, we train our model using both T1 and T2 modalities, which allows it to be fine-tuned on either T1 or T2 data alone during inference without requiring any modifications to the network structure. This is achieved by simply configuring the two input channels to process the same type of data (either T1 or T2).
Furthermore, if the downstream tasks involve more modalities than those used in the initial training stage on UK Biobank data, we can accommodate this by increasing the number of input channels. The pre-trained weights are then loaded into the corresponding layers of the network.
To investigate the efficacy of this method, we performed ablation testing on BraTS by restricting the modalities available to the model. The model was given only T1 or T2 images rather than utilizing all four standard modalities (T1, T2, T1ce, FLAIR) available in BraTS. The models were then trained with our Stage 2 pretraining and finetuning pipelines with these restrictions. Hyperparameters and number of GPUs were kept the same as in our earlier pretraining steps on BraTS as summarized in Table 3.
@@ 2.4.4 Modality Restriction and Flexibility of Modalities in Training and Inference
Since the ATLAS dataset has only one modality, T1-ce, the first layers of the Stage-1 pretrained model were modified by dropping the channel corresponding to the T2 modalitiy present in the UKB. For pretraining, four NVIDIA A100 GPUs, each with 32 GB of memory, were utilized. The stage 2 model took 35 hours to train, with a batch size and learning rate set to 4 and $5\times 10^{-3}$, respectively.
2.5.2 Fine-tuning on ATLAS v2.0
@@ 3.1 Pretraining
The pretraining of our BrainFounder models, which varied in size based on the number of parameters, took between 3 to 6 days. This process utilized a computational setup ranging from 8 to 64 NVIDIA A100 GPUs, each with 80GB capacity. Figure 4 illustrates the validation loss during the pretraining phase across different BrainFounder model sizes.
@@ 3.2.1 Comparison to State-of-the-Art Methods
Table 4 summarizes our BrainFounder’s best performing model against published results from other state-of-the-art models on the BraTS challenge. In addition, we include comparison to the corresponding single-stage pretrained model that was pretrained only on the UKB. The appendix present a comparative analysis of all trained BrainFounder models (of varying sizes) against the current leading model in this field broken down by model size.
Table 4: A comparison of BrainFounder models’ performance in terms of average Dice coefficient on the BraTS challenge. BF-S indicates our best performing BrainFounder mode. BF1-S indicates this model with only the Stage-One pertaining and without the Stage-Two pretraining step. SwinU is a comparison to the SwinUNETR model, with SwinU-MRI comparing to the model for BraTS published on GitHub, SwinU-Res containing the same modality restrictions as our current model, and SwinU-CT being pretrained using CT data.
BF-S (64M) BF1-S (64M) SwinU-MRI SwinU-Res SwinU-CT nn-UNET SegResNet TransBTS Model-Zoo Fold 1 0.9032 0.8994 0.8854 0.895 0.894 0.896 0.899 0.883 0.857 Fold 2 0.9182 0.9055 0.9059 0.899 0.902 0.917 0.916 0.902 0.879 Fold 3 0.9121 0.9125 0.8981 0.894 0.898 0.910 0.909 0.889 0.820 Fold 4 0.9100 0.9133 0.8924 0.890 0.893 0.909 0.908 0.893 0.889 Fold 5 0.9139 0.9114 0.9035 0.903 0.902 0.909 0.906 0.892 0.893 Average 0.9115 0.9110 0.8971 0.896 0.898 0.908 0.907 0.891 0.868
@@
  • Across all folds, the BrainFounder-Small framework consistently outperformed the SwinUNETR model. This indicates that the additional training steps taken within the BrainFounder framework play a significant role in enhancing its effectiveness in brain tumor segmentation tasks.
  • The Small (64M) parameter version of BrainFounder achieved higher Dice coefficients on average than the 62M and 69M versions. We believe this indicates that there is an optimal range of model complexity that maximizes performance, and simply increasing the number of parameters does not necessarily lead to better results.
  • The one-stage Tiny model performed comparably to the two-stage BrainFounder (62M) model, which is notable implying it did not benefit considerably from the second stage pretraining on the BraTS. This might imply that the UKB dataset alone provides enough variability for effective training. Further study should be made to verify whether the benefit of pretraining on the target datasets can be found using large-scale networks.

### v3

@@ 1.1.2 BRaTS dataset
In Stage 2 and Stage 3, one of the datasets we used to perform self-supervised pretraining and finetuning on MRI images is from the training set of the BraTS 2021 Task 1 (Tumor Segmentation) challenge. This dataset consists of 1,251 subjects each with T1w, T1-contrast enhanced (T1-ce), T2-weighted (T2w), and T2-FLAIR images. We obtained all publicly available imaging volumes as part of the challenge. The BraTS challenge utilizes a standard preprocessing pipeline similar to the UKB dataset. First, images are converted from DICOM images to NIfTI using dcm2niix ([li_first_2016]) and tools available from the Cancer Imaging Phenomics Toolkit (CaPTk) ([davatzikos_cancer_2018]). Images are then co-registered to the SRI24 template and resampled to a uniform resolution of 1 mm3. Finally, each modality is skull stripped, defaced, and converted to NIfTI ([baid_rsna-asnr-miccai_2021]). Imaging volumes were then segmented into three tumor classes using the STAPLE algorithm ([warfield_simultaneous_2004]) across previous BraTS winners and refined manually. These manual annotations were further verified by multiple board-certified neuro-radiologists, resulting in quality-controlled tumor segmentation labels across all four modalities in 3 classes: Gd-enhancing tumor (referred to as the whole tumor (WT), edematous tissue (ED), and necrotic tumor core (TC).
@@
Additionally, we perform self-supervised Stage 2 pretraining and fine-tuning on MRI images from the training set of the Anatomical Tracings of Lesions After Stroke (ATLAS) v2.0 Dataset [liew_large_2022]. This dataset consists of 655 T1-ce MRIs aggregated from 44 research cohorts. Each MRI is from one subject, and time points range from $<$24 hours to $>$180 days after stroke onset. The standard labeling pipeline for the ATLAS dataset consists of (1) manual quality control to exclude significant motion artifacts, (2) manual lesion tracing in ITK-SNAP [yushkevich_user-guided_2006] [yushkevich_itk-snap_2017], and (3) lesion mask review by two independent raters. The data then went through a similar prepossessing pipeline to BraTS. The MR images were intensity-normalized and registered to the MNI-152 template using the MINC toolkit (https://github.com/BIC-MNI/minc-toolkit). Finally, FreeSurfer’s MRI deface functionality was used to deface the scans. Images were reviewed again in a final quality check at the end of the pipeline before being included in the dataset. Segmentations are evaluated on four metrics - Dice coefficient for the final segmentation (Dice), the difference between true total lesion volume and predicted total lesion volume (Volume Difference), the difference in the number of lesions between ground truth and prediction (Lesion Count), and Lesion-wise F1 Score. The Lesion-wise F1 Score is calculated by performing a 3D connected-component analysis to determine true positives, false positives, and false negatives. A true positive is any 3D connected component in the ground-truth image that overlaps with at least one voxel in the prediction image. Conversely, a false positive is any 3D connected component in the prediction image that does not overlap with the ground-truth image. A false negative is a connected component in the ground truth that lacks overlapping voxels in the prediction image [liew_large_2022].
@@ 1.2 Stage 1: Pretraining on the UKB
The initial pretraining stage involves the self-supervised learning of a transformer-based neural network model using a substantial unlabeled image dataset. For this purpose, the UKB dataset ([littlejohns_uk_2020]) is utilized. From our 82,800 3D volumetric images used for pretraining, the input MRI modalities are randomly cropped into $96\times 96\times 96$ sub-volumes and augmented with random inner cutout and rotation. These augmented images are then fed into the SwinUNETR encoder for processing.
The SwinUNETR architecture incorporates a Swin Transformer encoder that handles 3D input patches. This encoder operates with a patch size of $2\times 2\times 2$, a feature dimension of 8, and an embedding space of 48 dimensions. It consists of four stages, with a patch merging layer introduced between stages to reduce the feature size by half.
Adopting the methodology from [tang_self-supervised_2022], the SwinUNETR encoder is pretrained through three distinct proxy tasks that serve as self-supervised fine-tuning mechanisms: masked volume inpainting, 3D image rotation, and contrastive coding. The primary objective of pretraining is to minimize the total loss function. This work has developed three models to address varying complexities. These models include the foundational BrainSegFounder-Tiny with 62 million parameters, the intermediate BrainSegFounder-Small with 64 million parameters, and the advanced BrainSegFounder-Big with 69 million parameters. The primary differentiation among these models is the variation in the number of sliding window blocks within their third stage. Table 2 shows BrainSegFounder’s sliding-window encoder backbone’s parameters, number of SSL heads, and number of sliding window blocks.
For the pretraining process, 64 NVIDIA DGX A100 GPUs, distributed across 8 DGX-2 nodes, are deployed at the University of Florida’s HiPerGator-AI supercomputer. Data parallelism is implemented to optimize the efficiency of model training. Both training and validation losses are monitored to track progress. The AdamW optimizer is employed using a warm-up cosine scheduler set for 500 iterations. The training employs a batch size of 2 per GPU, using $96\times 96\times 96$ patches. The initial learning rate is established at $6\times 10^{-6}$, coupled with a momentum of 0.9 and a decay of 0.1 over 15,000 iterations. These parameters are summarized in Table 3.
BrainSegFounder-Tiny (62M) BrainSegFounder-Small (64M) BrainSegFounder-Big (69M)
# of encoder parameters 19,097,191 20,982,103 26,636,839
Encoder layer level Output size # of SSL Heads # Swin Blocks # of SSL Heads # Swin Blocks # of SSL Heads # Swin Blocks
Level 1 48x(48x48x48) 3 2 3 2 3 2
Level 2 96x(24x24x24) 6 2 6 2 6 2
Level 3 192x(12x12x12) 12 2 12 6 12 18
Level 4 384x(6x6x6) 24 2 24 2 24 2
Feature size 48 48 48
Bottleneck dimension 768 768 768
@@
Fig. 4: Training (left) and validation (right) loss of Stage 1-pretraining three different scale of BrainSegFounder models on UKB.
@@ 1.3.1 Stage 2: Pretraining on BraTS
Given that the BraTS dataset comprises four modalities, but only two (specifically, T1w and T2-FLAIR) were available for pretraining in the initial stage, the first layer of the pretrained network on UKB was modified. This modification involved expanding the number of input channels by adding two new channels, whose weights were randomly initialized using the Kaiming initialization method ([he_delving_2015]).
Hyperparameter settings for Stage-2 Pretraining can be found in Table 3. For pretraining on BraTS, two NVIDIA A100 GPUs, each with 32 GB of memory, were utilized. Depending on the model size, the BrainSegFounder models require between 48 to 72 hours for training. The batch size and learning rate were uniformly set at 2 and $1\times 10^{-4}$ for all models during this pretraining phase.
1.3.2 Stage 3: Fine-tuning on BraTS
@@ 1.3.3 Few-shot Learning on BraTS
To investigate our model’s performance using limited training data, we conducted a systematic comparison between BrainSegFounder and the baseline model, SwinUNETR, utilizing a descending percentage training approach in the context of the BraTS challenge. Using both our BrainSegFounder pretrained model and SwinUNETR, we finetuned on 40% of the BraTS training dataset, with subsequent incremental reductions in data availability, decreasing to a final 5% of the original dataset. Due to the potential high-variability on the input dataset at such small percentages of input data, we trained on 5 different randomly sampled subsets of the input training data and calculated the average performance across these subsets. This method aimed to explore the impact of training data scarcity on model performance and adaptability. Performance evaluations were carried out on the BraTS test set after each training step and evaluated with the Dice coefficient to assess segmentation accuracy.
@@
Stage Data No. Subjects GPU Batch size Learning rate No. steps No. input channel
Encoder only Stage 1 Pretraining UKB 43369 64 $\times$ A100 128 $1\times 10^{-6}$ $2\times 10^{5}$ 2
Encoder + Decoder Stage 2 Pretraining & Stage 3Fine-tuning BRaTS 1251 2 $\times$ A100 2 $1\times 10^{-4}$ 50000 4
Encoder + Decoder Stage 2 Pretraining & Stage 3 Fine-tuning ATLAS v2.0 655 4 $\times$ A100 4 $3\times 10^{-3}$ 600 1
1.3.4 Modality Restriction and Flexibility in Training and Inference
The proposed method is designed to be adaptable across various data modalities in downstream tasks. In the case that fewer modalities are available in the downstream task, Stage 1 pretrained model using both T1- and T2-weighted MRI can be adapted and fine-tuned on fewer modalities (e.g., T1- or T2- weighted MRI alone) during Stage 2 pretraining, Stage 3 supervised training, and the inference without requiring any modifications to the network structure. This is achieved by simply configuring the two input channels to process the same type of data (either T1- or T2-weighted).
In the case that more modalities are available in the downstream task, our model can also accommodate this by increasing the number of input channels. The pre-trained weights are then loaded into the corresponding layers of the network.
To investigate the efficacy of this method, we performed ablation testing on BraTS by restricting the modalities available to the model. The Stage 1 model was given only either T1w or T2w images rather than utilizing both modalities available from the UKBiobank. These models, pretrained on only one modality, were then trained with our Stage 2 pretraining pipeline on the BraTS data with only the modality trained on in Stage 1 instead of the 4 modalities available from BraTS. Finally, the model was finetuned on BraTS with the same single modality. Hyperparameters and number of GPUs were kept the same as in our earlier pretraining steps on BraTS as summarized in Table 3.
@@ 1.3.4 Modality Restriction and Flexibility in Training and Inference
Since the ATLAS dataset has only one modality, T1-ce, the first layers of the Stage-1 pretrained model were modified by dropping the channel corresponding to the T2w modalitiy present in the UKB. For pretraining, four NVIDIA A100 GPUs, each with 32 GB of memory, were utilized. The stage 2 model took 35 hours to train, with a batch size and learning rate set to 4 and $5\times 10^{-3}$, respectively.
1.4.2 Stage 3: Fine-tuning on ATLAS v2.0
@@ 2.1 Pretraining
The pretraining of our BrainSegFounder models, which varied in size based on the number of parameters, took between 3 to 6 days. This process utilized a computational setup ranging from 8 to 64 NVIDIA A100 GPUs, each with 80 GB capacity. Figure 4 illustrates the validation loss during the pretraining phase across different BrainSegFounder model sizes.
@@ 2.2.1 Comparison to State-of-the-Art Methods
Table 4 summarizes our BrainSegFounder’s best performing model against published results on our validation splits from other state-of-the-art models on the BraTS challenge: the baseline SwinUNETR model [hatamizadeh_swin_2022], nnU-Net [isensee_nnu-net_2021], TransBTS [wang_transbts_2021], SegResNet [myronenko_3d_2018_segresnet], and MONAI’s Model-Zoo [project-monaimodel-zoo_2024]. SegResNet and nnU-Net are both winning methodologies in previous BraTS challenges. TransBTS is a vision-transformer based approach tailored for brain tumor segmentation. MONAI’s Model-Zoo is a bundle of medical imaging models capable of performing a wide variety of tasks, including BraTS segmentation. In addition, we include comparison to the corresponding single-stage pretrained model that was pretrained only on the UKB. Table 8 in the appendix present a comparative analysis of all trained BrainSegFounder models (of varying sizes) against the current leading model in this field broken down by model size.
Table 4: A comparison of BrainSegFounder (BSF) models’ performance in terms of average Dice coefficient on the BraTS challenge. BSF-S indicates our best performing BrainSegFounder model (small, 64M parameters). BrainSegFounder models were pretrained with SSL on T1- and T2-weighted MRI 3D volumes and finetuned with supervised learning using all four modalities present in BraTS. BSF-1-S indicates this model with only the Stage 1 (SSL) pertaining on UKB and without the Stage 2 pretraining step. SwinU models are models using the SwinUNETR architecture trained on BraTS via supervised learning. SwinU-MRI is the model trained directly using supervised learning on BraTS published on GitHub (https://github.com/Project-MONAI/research-contributions/tree/main/SwinUNETR/BRATS21), SwinU-Res is pretrained with SSL on only T1w and T2w and finetuned on BraTS, and SwinU-CT pretrained using CT data and finetuned with supervised learning on BraTS. nnU-Net and SegResNet are former BraTS challenge winners trained using supervised learning on our folds. TransBTS is a vision-transformer based segmentation algorithm optimized for brain segmentation. Model-Zoo is a bundle of models published by MONAI that can perform BraTS segmentation out of the box using their "Brats mri segmentation" [sic] model found at https://monai.io/model-zoo.html
BSF-S (64M) BSF-1-S (64M) SwinU-MRI SwinU-Res SwinU-CT nnU-NeT SegResNet TransBTS Model-Zoo
Fold 1 0.9032 0.8994 0.8854 0.895 0.894 0.896 0.899 0.883 0.857
Fold 2 0.9182 0.9055 0.9059 0.899 0.902 0.917 0.916 0.902 0.879
Fold 3 0.9121 0.9125 0.8981 0.894 0.898 0.910 0.909 0.889 0.820
Fold 4 0.9100 0.9133 0.8924 0.890 0.893 0.909 0.908 0.893 0.889
Fold 5 0.9139 0.9114 0.9035 0.903 0.902 0.909 0.906 0.892 0.893
Average 0.9115 0.9110 0.8971 0.896 0.898 0.908 0.907 0.891 0.868
@@
  • Across all folds, the BrainSegFounder-Small (BSF-S) framework consistently outperformed the SwinUNETR model. This indicates that the additional training steps taken within the BrainSegFounder framework play a significant role in enhancing its effectiveness in brain tumor segmentation tasks.
  • The Small (64M parameters) version of BrainSegFounder achieved higher Dice coefficients on average than the 62M and 69M versions. We believe this indicates that there is an optimal range of model complexity that maximizes performance given certain training data size. Simply increasing the number of parameters does not necessarily lead to better results if the training data does not scale up.
  • The one-stage Tiny (62M parameters) model performed comparably to the two-stage BrainSegFounder-Tiny (62M parameters) model, which is notable, implying it did not benefit considerably from the second stage pretraining on the BraTS. This might imply that the UKB dataset alone provides enough variability for effective training. Further study should be made to verify whether the benefit of pretraining on the target datasets can be found using large-scale networks.